# Moto: Latent Motion Token as the Bridging Language for Robot Manipulation

### v1

@@ 4.1 Benchmarks and Datasets
On the SIMPLER benchmark, we focus on three tasks concerning the Google Everyday Robot embodiment: Pick Coke Can, Move Near, and Open/Close Drawer, as illustrated in Figure 4. The “Pick Coke Can” task involves grasping and lifting the empty coke can in three different orientations: horizontal laying, vertical laying, and standing. The “Move Near” task places 3 (out of 8) objects in a triangle pattern on the tabletop and instructs the robot to move a designated source object near another object as the target. We utilize a subset of Open-X-Embodiment [52] to train the Latent Motion Tokenizer and pre-train Moto-GPT, which consists of 109k real-world trajectory videos [5, 53, 49, 41, 14, 37, 39, 45, 61, 10, 60] across various embodiments. For fine-tuning Moto-GPT, we use 73k action-labeled expert trajectories from the RT-1 Robot Action dataset [5].
@@ 4.2 Compared Models
On the SIMPLER benchmark, we compare Moto-GPT with four representative models pre-trained with Open-X-Embodiment datasets:
@@
  • RT-2-X [62] adapts the pre-trained large vision-language model (VLM), PaLI-X (55B) [11], into a robot policy by casting tokenized actions into text tokens.
@@
  • OpenVLA [28] builds on a pre-trained Prismatic-7B [26] VLM backbone for robot action prediction.
@@
  • RoboFlamingo [32] is a robot policy model adapted from OpenFlamingo [2], a large VLM pre-trained on extensive vision-language corpus.
@@
We also study the following variations of Moto-GPT as optional baselines:
@@
4.3 Training Details
Latent Motion Tokenizer.
The implementation details for the trainable modules of the Latent Motion Tokenizer are summarized in Table 1. We use the hyperparameters listed in Table 2 to train this model on four A100-40G GPUs. To facilitate the learning of latent motion tokens, we downsample the original videos in the training dataset, ensuring that the visual motion between frames is sufficiently distinct. Specifically, for videos from the Open-X-Embodiment datasets, we sample one frame every three frames (i.e., $\Delta t=3$) and train the Latent Motion Tokenizer for 350k steps. For videos from the CALVIN dataset, we adopt a sampling rate of one frame every five frames ($\Delta t=5$) and train the model for 150k steps.
@@ 4.3 Training Details
Moto-GPT.
We present the implementation details of Moto-GPT in Table 3, where the Action Head is included only during the fine-tuning phase. Moto-GPT handles a maximum video length of three frames, and the video downsampling rate applied during both the pre-training and fine-tuning stages is consistent with the rate used for training the Latent Motion Tokenizer. When fine-tuning Moto-GPT across different benchmarks, the number of action query tokens inserted after the latent motion tokens at each time step varies. Specifically, for the SIMPLER benchmark, we insert three action query tokens, whereas for the CALVIN benchmark, we insert five. For pre-training, Moto-GPT is trained for 10 epochs using eight A100-40G GPUs, with the relevant hyperparameters outlined in Table 4. The hyperparameters for fine-tuning remain consistent with those used during pre-training, except for the number of epochs. We fine-tune Moto-GPT for three epochs on the RT1-Robot-Action dataset and 18 epochs on the CALVIN dataset, utilizing four A100-40G GPUs.
5 Experiments
5.1 Latent Motion Token as an Interpretable Motion Language
As illustrated in Figure 6, the next frame reconstructed by the Latent Motion Tokenizer using ground-truth latent motion tokens is authentic, effectively capturing the key dynamics between the initial frame and the ground-truth next frame. This suggests that latent motion tokens can represent fine-grained motion details, and the Latent Motion Tokenizer’s decoder serves as a qualified simulator for visualizing environmental changes.
Figure 7 further explores the controllability and consistency of latent motion tokens. Each row demonstrates that different token chunks produce visual motions with varying orientations and scales relative to the initial frame. Conversely, within each column, identical token chunks yield similar effects on the resulting positions and postures across different starting observations. By concatenating latent motion token chunks for every two consecutive frames from a video, we create a sequential representation of motion trajectories, akin to natural language context. As shown in Figure 8, this representation can be applied to different initial observations, generating contextualized motion trajectories and highlighting the potential of latent motion tokens as a unified language interface for guiding imitation learning.
[…]
@@ 5.1 Latent Motion Token as an Interpretable Motion Language
Figure 7: Visualization of latent motion token interpretability. Each row displays reconstructed frames from the same initial frame using different latent motion tokens, while each column shows frames reconstructed from the same latent motion tokens with varying initial frames. The latent motion tokens exhibit consistent (see columns) and discriminative (see rows) semantics, despite being trained in an unsupervised manner.
Figure 8: Video imitation generation via latent motion tokens, where a sequence of latent motion tokens from a demonstration video are extracted by the Latent Motion Tokenizer and are decoded into a new video. This generated video is based on a different initial frame while preserving the original robot movement semantics.
5.2 Moto-GPT as a Useful Motion Prior Learner
The pre-training stage of Moto-GPT involves autoregression on video data using latent motion tokens, enabling it to predict motion trajectories based on initial observations and various language prompts, as illustrated in Figure 9. Table 6 presents the top-k accuracy of Moto-GPT in predicting ground-truth latent motion tokens from a 128-size codebook on the validation splits of the pre-training datasets. These results demonstrate Moto-GPT’s effective acquisition of prior knowledge for motion trajectory prediction, which is crucial for robot action inference based on human instructions. Thus, the learned motion priors hold the potential to benefit downstream robotic tasks.
Video Representation Semantic Acc.
Initial frame 0.292
Initial frame repeated by 8 times 0.283
Initial frame + 7 subsequent frames 0.828
Initial frame + 7 latent motion token chunks 0.797
Additionally, latent motion tokens allow Moto-GPT to interpret trajectory videos as compact token sequences and evaluate their rationality through the autoregressive likelihood defined in Eq. 1. Figure 10 illustrates the potential of using Moto’s log-likelihoods as a reward signal for trajectory videos, indicating how well a trajectory aligns with Moto-GPT’s distribution and measuring the temporal consistency of behavior. To assess this, we collected 98 video triplets in CALVIN using the baseline policies and a random policy. Each triplet consists of three types of trajectory videos originating from the same environment state. The averaged log-likelihoods for each trajectory type at each sequence step, shown in Figure 10, clearly differentiate successful trajectories from failures and random attempts.
5.3 Moto-GPT as an Effective Robot Policy
Table 6: Top-K motion token prediction accuracy of Moto-GPT.
@@ 5.3 Moto-GPT as an Effective Robot Policy
Method Pick Coke Can Move Near Open / Close Drawer Overall
RT-1-X [5] 0.820 0.330 0.550 0.567 0.317 0.296 0.891 0.597 0.534
RT-2-X [62] 0.740 0.740 0.880 0.787 0.779 0.157 0.343 0.250 0.607
Octo-Base [42] 0.210 0.210 0.090 0.170 0.042 0.009 0.444 0.227 0.169
OpenVLA [28] 0.270 0.030 0.190 0.163 0.462 0.194 0.518 0.356 0.248
Moto 0.820 0.500 0.900 0.740 0.604 0.130 0.732 0.431 0.614
[…]
Model Observation Space Tasks competed in a row (1000 chains)
1 2 3 4 5 Avg. Len.
SuSIE [4] Static RGB 0.870 0.690 0.490 0.380 0.260 2.69
Moto Static RGB 0.897 0.729 0.601 0.484 0.386 3.10
Moto w/o Motion Token Static RGB 0.779 0.555 0.380 0.256 0.167 2.14
Overall Performance.
After fine-tuning, we evaluated Moto-GPT [^3] For simplicity, we will refer to Moto-GPT as Moto in the following experimental tables and figures. against baseline models on the SIMPLER and CALVIN benchmarks, as shown in Tables 7 and 8. Overall, Moto-GPT outperforms the baselines on both benchmarks. Notably, on SIMPLER, Moto-GPT surpasses larger models like RT-2-X (PaLI-X 55B) and OpenVLA (Prismatic-7B), despite having only 98M parameters for the GPT-style backbone. Moto-GPT also shows strong generalization in the unseen CALVIN environment. The baseline models utilize various pre-training strategies: SuSIE employs a pre-trained image-editing model for goal image generation, RobotFlamingo is initialized from a large vision-language model, MT-R3M uses a pre-trained robot visual encoder, and GR-1 predicts future pixel values based on input observations. In contrast, Moto-GPT, pre-trained through autoregressive motion token prediction, achieves competitive performance despite relying solely on RGB images from a static camera. This is particularly impressive when compared to GR-1, which uses images from both static and gripper cameras along with proprioceptive robot state data. Our findings support the idea that focusing on motion-related dynamics rather than frame-level visual details is a more effective approach for learning from videos. Additionally, Moto-GPT significantly outperforms its variant trained from scratch on action-labeled robot data without latent motion tokens (Moto w/o Motion Token). This highlights the effectiveness of our latent-motion-token-based pre-training and co-fine-tuning strategy in enhancing policy performance for practical robot manipulation tasks.
Data Efficiency.
[…]
Ablations on Policy Fine-tuning Methods.
In Figure 12, we evaluate the effectiveness of Moto’s co-fine-tuning strategy. Moto-IML and Moto-DM share the same pre-training approach as Moto-GPT but differ in their fine-tuning methods. Specifically, Moto-IML omits the loss term for latent motion token prediction, while Moto-DM excludes latent motion tokens from the input sequence entirely. When compared to Moto w/o Motion Tokens, which is trained from scratch without latent motion tokens, both Moto-IML and Moto-DM show performance improvements due to the motion priors gained during pre-training. However, they still fall short of Moto-GPT’s performance. This highlights the importance of retaining latent motion tokens in the sequence, allowing action query tokens to transfer knowledge through direct attention. Furthermore, co-fine-tuning for latent motion token prediction helps preserve the learned motion priors in Moto-GPT.
[…]
6 Conclusion and Discussion
@@
Firstly, Moto demonstrates the feasibility of learning a unified language to interpret diverse visual dynamics from videos, eliminating the need for hardware-specific action labels. The latent motion trajectories tokenized from videos provide a rich resource for models to learn motion priors closely related to low-level actions. While we currently mainly use robot videos to train the Latent Motion Tokenizer, the learned latent motion tokens demonstrate the potential to produce consistent visual motions across varied contexts and embodiments. We believe a similar approach could be applied to human motion representation, enabling models to learn a wealth of world knowledge from Internet-scale human videos.
Besides, the Moto-GPT pre-trained on videos tokenized into latent motion token sequences and fine-tuned on action-labeled trajectories, effectively transfers motion priors learned from videos to actual robot action prediction. This is particularly beneficial in low-resource scenarios. Future work could involve scaling up pre-training video data and optimizing fine-tuning to improve model performance on downstream robot tasks further.

### v4

@@ 7.1 Benchmarks
On the SIMPLER benchmark, we focus on three tasks concerning the Google Everyday Robot embodiment: Pick Coke Can, Move Near, and Open/Close Drawer, as illustrated in Fig. 13. The “Pick Coke Can” task involves grasping and lifting the empty coke can in three different orientations: horizontal laying, vertical laying, and standing. The “Move Near” task places 3 (out of 8) objects in a triangle pattern on the tabletop and instructs the robot to move a designated source object near another object as the target. The “Open/Close Drawer” task requires the robot to open or close a specific drawer from the top, middle, or down position of a cabinet.
@@
Real-world Robot Experiments.
We design three tasks for real-world evaluation: Pick-Place Banana (picking up a banana from the tabletop and placing it in a pan), Close Laptop (pushing the laptop’s lid until it is completely closed), and Disassembly (removing the stick that is assembled in the slot). All tasks are executed with a FANUC LR Mate 200iD robot arm, as shown in Fig. 3. Each task is tested for 10 times, with the initial positions of objects randomized. For generalization evaluation, we consider two scenarios: (i) Novel Object: we change the color, texture, and shape of the manipulated object; (ii) Visual Distractor: we add irrelevant objects as distractors.
7.2 Datasets
Training Latent Motion Tokenizer & Pre-training Moto-GPT.
On the SIMPLER benchmark, we use a subset of Open-X-Embodiment (OXE) [52] to train the Latent Motion Tokenizer and pre-train Moto-GPT, including 109k real-world trajectory videos from RT-1 Robot Action [4], Bridge [53], Task Agnostic Robot Play [48, 40], Jaco Play [13], Cable Routing [36], RoboTurk [38], NYU VINN [44], Austin VIOLA [61], Berkeley Autolab UR5 [9], and TOTO [60] datasets across various embodiments. On CALVIN, we use all the play videos from environments A, B, and C to train the Latent Motion Tokenizer. 35% data (including 18k trajectory videos) with language annotations is used for autoregressive pre-training. The real-world experiments also use OXE data for pre-training.
Fine-tuning Moto-GPT.
On the SIMPLER benchmark, we use the 73k action-labeled real-world expert trajectories from RT-1 Robot Action [4] to fine-tune the policy model. On the CALVIN benchmark, we use the 18k demonstration trajectories with language annotations and action labels from environments A, B, and C for fine-tuning. For real-world experiments, we collect 90 demonstration trajectories (30 for each task) with teleoperation for fine-tuning.
Note that all the pre-training and fine-tuning data for SIMPLER come from the real world instead of simulation environments. This setting aims to study the model’s transfer ability between real and simulation scenarios. On the other hand, the ABC-only setting for CALVIN training data aims to evaluate the model’s zero-shot generalization capability to the unseen environment D.
@@ 7.3 Compared Models
On the SIMPLER benchmark, we mainly compare Moto-GPT with five representative models pre-trained with Open-X-Embodiment datasets:
@@
  • RT-2-X [62] adapts the pre-trained large vision-language model (VLM), PaLI-X (55B), into a robot policy by casting tokenized actions into text tokens.
@@
  • OpenVLA [27] builds on a pre-trained Prismatic-7B VLM backbone for robot action prediction.
  • OpenVLA (fine-tuned) further fine-tunes OpenVLA on the RT-1 Robot Action dataset [4], despite its action-labeled pre-training data already contains this dataset.
@@
  • RoboFlamingo [32] is a robot policy model adapted from OpenFlamingo, a large VLM pre-trained on extensive vision-language corpus.
@@
We study the following variations of Moto-GPT as optional baselines:
@@
8 Training Details
8.1 Latent Motion Tokenizer
The implementation details for the trainable modules of the Latent Motion Tokenizer are summarized in Table 4. We use the hyperparameters listed in Table 5 to train this model on four 40G GPUs. To facilitate the learning of latent motion tokens, we downsample the original videos in the training dataset, ensuring that the visual motion between frames is sufficiently distinct. Specifically, for videos from the OXE data, we sample one frame every three frames (i.e., $\Delta t=3$) for videos from the RT-1 Robot-Action subset. The sampling rates for videos from other OXE subsets are adjusted based on the ratio of their fps to that of RT-1 Robot-Action videos. For human videos, $\Delta t$ is empirically set to 6. We train the Latent Motion Tokenizer for 350k steps. For videos from the CALVIN dataset, we adopt a sampling rate of one frame every five frames ($\Delta t=5$) and train the model for 150k steps. For real-world robot experiments, we fine-tune the Latent Motion Tokenizer pre-trained on OXE videos for another 500 steps on the newly collected real-world trajectory videos.
@@ 8.1 Latent Motion Tokenizer
8.2 Moto-GPT
We present the implementation details of Moto-GPT in Table 6, where the Action Head is included only during the fine-tuning phase. Moto-GPT handles a maximum video length of three frames, and the video downsampling rate applied during both the pre-training and fine-tuning stages is consistent with the rate used for training the Latent Motion Tokenizer. When fine-tuning Moto-GPT across different benchmarks, the number of action query tokens inserted after the latent motion tokens at each time step varies. Specifically, for the SIMPLER benchmark, we insert three action query tokens, whereas for the CALVIN benchmark, we insert five. For pre-training, Moto-GPT is trained for 10 epochs using eight 40G GPUs, with the relevant hyperparameters outlined in Table 7. The hyperparameters for fine-tuning remain consistent with those used during pre-training, with the exception of the number of epochs. During fine-tuning, Moto-GPT is trained for three epochs on the RT1-Robot-Action dataset and 18 epochs on the CALVIN dataset, utilizing four 40G GPUs. For real-world experiments, we start with the same pre-trained checkpoint of Moto-GPT as adopted for the SIMPLER benchmark. We further pre-train it with a combination of OXE videos and the 90 newly collected trajectory videos for 5 epochs. Then we fine-tune it with real robot action labels for 20 epochs.
9 Details of Experiments
@@
Figure 16: Predicted video trajectories by the pre-trained Moto-GPT for CALVIN tasks reflecting delicate robot actions.
Table 8: Top-K motion token prediction accuracy of Moto-GPT in predicting ground-truth latent motion tokens from a 128-size codebook on the validation splits of the pre-training datasets.
@@
10 Limitations & Future Work
@@
Firstly, Moto demonstrates the feasibility of learning a unified language to interpret diverse visual dynamics from videos, eliminating the need for hardware-specific action labels. The latent motion trajectories tokenized from videos provide a rich resource for models to learn motion priors closely related to low-level actions. While we currently mainly use robot videos to train the Latent Motion Tokenizer, the learned latent motion tokens demonstrate the potential to produce consistent visual motions across varied contexts and embodiments. Notably, our preliminary experiments with SSV2 videos show promising results in human-to-robot motion transfer via latent motion tokens. We believe a similar approach could be applied to model more complex human motions, enabling robots to learn a wealth of world knowledge from Internet-scale videos.
Besides, the Moto-GPT pre-trained on videos tokenized into latent motion token sequences and fine-tuned on action-labeled trajectories, effectively transfers motion priors learned from (even human) videos to actual robot action prediction. This is particularly beneficial in low-resource scenarios. Future work could scale up pre-training video data and optimize fine-tuning to improve model performance on downstream robot tasks further.